OpenStudy (anonymous):

im stuck on a question can someone help me out? 2.Write each as a power of 2, show your work. A.computer capacity is often measured in bits and bytes. A bit is a smaller unit, which is 1 or 0 in the computers memory. a byte is 2^3 bits. a megabyte (MG) is 2^20 bytes. How many bits are in a megabyte? B. A gigabyte (GB) is 2^30 bytes. How many megabytes are in a gigabyte? How many bits are in a gigabyte?

OpenStudy (anonymous):

2^3 and 2^20 POWER for a you add or times and B B you 2^30 power you do 2 and 30 times or divide.

OpenStudy (anonymous):

i think thats what you do if i can rember the order right

OpenStudy (lcnettech):

a 20*2=40 and a megabyte is 40 b 20*3=60 and their are 60 gigabytes

OpenStudy (anonymous):

thanks ^.^ @lcnettech

OpenStudy (lcnettech):

your welcome

OpenStudy (lcnettech):

anything else

OpenStudy (anonymous):

whats B?

OpenStudy (lcnettech):

20*3=60 and a gigabyte is 60

OpenStudy (anonymous):

ok thanks

OpenStudy (lcnettech):

your welcome